OpenStudy (anonymous):

help!! A regular octagon is inscribed in a circle of radius 11.0 centimeters. Approximate the perimeter of the octagon. (Round your answer to the nearest tenth.)

OpenStudy (anonymous):

Perimeter = 8 * sides = 8 * 2r = 8 * 22 = 176 cm